OpenStudy (anonymous):

Which of the following is the general solution of the differential equation dy/ dx=2x/y a: y^2 = x^2 + C b: y^2 = 2x^2 + C c: y^2 = 4x^2 + C d: x^2 − y^2 = C

OpenStudy (anonymous):

@nirmalnema can u plz help? :)

OpenStudy (anonymous):

take dy/dx for a,b,c,d and see which is one look like dy/dx=2x/y

OpenStudy (nirmalnema):

dy/ dx=2x/y ydy=2xdx upon int. y^2/2=x^2+c y^2=2x^2+c

OpenStudy (anonymous):

a) dy/dx y^2=x^2+c 2y dy/dx=2x .... => dy/dx=2x/2y=x/y

OpenStudy (anonymous):

so is not the answer

OpenStudy (anonymous):

so it is b

OpenStudy (anonymous):

you guys both give great answers i dont know who to medal!!!

OpenStudy (anonymous):

thank u :)